OpenStudy (anonymous):

Help Ill give a medal who answers this! The change in water level of a lake is modeled by a polynomial function, W(x). Describe how to find the x-intercepts of W(x) and how to construct a rough graph of W(x) so that the Parks Department can predict when there will be no change in the water level. You may create a sample polynomial to be used in your explanations.

OpenStudy (campbell_st):

to find the x intercepts, let W(x) = 0 and solve for x here is a really simple example... let \[W(x) = -x^2 - 5x + 6\] let W(x) = 0 so \[0 = x^2 + 5x - 6\] this can be solved by factoring -(x -3)(x-2) = 0 so x = 2 and x = 3 are the intercepts... and the graph looks like